OpenStudy (abmon98):

T=T0e^-0.5t show that t=2lnT0/T Hence find t when T0=30 and T=10

OpenStudy (abmon98):

\[T=T _{0}e^-0.5t\] \[2\ln*(T _{0}/T)\]

OpenStudy (kropot72):

Do you know how to take In of both sides of the equation? That is the first step.

OpenStudy (abmon98):

\[\ln T=(-0.5t)LnT _{0}\]

OpenStudy (kropot72):

Not quite. \[\ln (ab)=\ln a+\ln b\] Therefore: \[\ln T=\ln T _{0}-0.5t\] Note: \[\ln e ^{-0.5t}=-0.5t\]

OpenStudy (kropot72):

The original equation can be written as follows: \[T=T _{0} \times e ^{-0.5t}\ ..........(1)\] Taking natural logs of both sides of (1) we get: \[\ln T=\ln T _{0}+\ln e ^{-0.5t}\ .........(2)\] But\[\ln e ^{-0.5t}=-0.5t\] Therefore we can rewrite (2) as: \[\ln T=\ln T _{0}-0.5t\ .......(3)\]

OpenStudy (kropot72):

@Abmon98 Does that explanation help you to understand?

OpenStudy (abmon98):

yes

OpenStudy (kropot72):

Good! Equation (3) can be rearranged as follows: \[0.5t=\ln T _{0}-\ln T\ .........(4)\] Can you make the right hand side of (4) a single Ln ?

OpenStudy (kropot72):

Hint: \[\ln a-\ln b=\ln \frac{a}{b}\]

OpenStudy (abmon98):

oh now i get it thank you so much

OpenStudy (kropot72):

You're welcome :)